Supplemental Operations Report – April 6th

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Power system monitor

Action taken for non-nominal systems: Still figuring out the best option for it. It needs to be hardwired into the internet, but also close enough to the ScienceDome to connect to the data

Generator: Still being manually run at night.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— Nominal

Notes on power system: The power system is acting crazy again. Yesterday we lost power due to another overload of the solar system. Then today the SOC is reading incorrectly again.I will look at it between crews.

Propane Reading, station tank – 60%

Propane Reading, director tank— 60%

Propane Reading, intern tank— 70%

Propane Reading, generator— 32 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 250 gallons

Sojourner rover used: no

Hours: 167.5

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Still haven’t started on the battery cleaning and watering.

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: 350.1 is blocking the gap between the campus and the outpost so no one will damage the repaired wiring conduit.

HabCar used and why, where? Yes, to town for garbage and water runs

CrewCar used and why, where? Yes, to town

Luna used and why, where? Yes, to town

Campus-wide inspection, if action is taken, what and why? Nothing to report.

General notes and comments: Nothing to report

Summary of general operations: Nothing to report

Summary of internet: We removed the modems and routers for the two HughesNet systems that have been deactivated in the Hab. Starlink will now be the only internet available on campus for crews. We do need to secure it with a password still.

Summary of suits and radios: Nothing to report

Summary of Hab operations: Evidence of rodents noted.

Summary of Outpost operations: Evidence of rodents noted.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Went into the robotic observatory today and restarted the system with Peter. Also realigned the small telescope’s camera which had become loose where the cords connect.

Summary of health and safety issues: Nothing to report.